OpenStudy (anonymous):

Whats the slope of the line through points (6,12) and (-6,2)

OpenStudy (mathstudent55):

slope = m = (difference in y values)/(difference in x values)

OpenStudy (irishboy123):

slope = dy/dx dy = 2-12 = -10 dx = -6-6 = -12 slope =5/6

OpenStudy (ajprincess):

Slope m of a line that passes through the points \((x_1, y_1)\) and \((x_2.y_2)\) can be found using the formula \[m=\frac{ y_2-y_1 }{ x_2-x_1}\] Does that help?

OpenStudy (anonymous):

The Answer has to be either -6/7,-7/6, 6/7 or 7/6

OpenStudy (ajprincess):

is the first point (6,-12)?

OpenStudy (anonymous):

no (6,12)

OpenStudy (ajprincess):

how about the second point?

OpenStudy (anonymous):

(-6,2)

OpenStudy (ajprincess):

if these are the points I am afraid the options given are wrong.:(

OpenStudy (anonymous):

okay here i go,we have a formula for finding slope given two points (x_1,y_1) and (x_2,y_2) m=y_1-y_2/x_1-x_2 or,m=12-2/6+6=10/12=5/6

OpenStudy (anonymous):

RIght? thats what i was thinkin its super confusing!:( and i got 5/6 too but that answer is not an option!://

OpenStudy (anonymous):

:/

OpenStudy (anonymous):

well do we know for sure if its ganna be a positive or negative? cuz then that will give me a 50/50 chance of getting it right..